# MambaXCTrack: Mamba-based Tracker with SSM Cross-correlation and Motion Prompt for Ultrasound Needle Tracking

### v1

@@ Abstract
Ultrasound (US)-guided needle insertion is widely employed in percutaneous interventions. However, providing feedback on the needle tip position via US image presents challenges due to noise, artifacts, and the thin imaging plane of US, which degrades needle features and leads to intermittent tip visibility. In this paper, a Mamba-based US needle tracker MambaXCTrack utilizing structured state space models cross-correlation (SSMX-Corr) and implicit motion prompt is proposed, which is the first application of Mamba in US needle tracking. The SSMX-Corr enhances cross-correlation by long-range modeling and global searching of distant semantic features between template and search maps, benefiting the tracking under noise and artifacts by implicitly learning potential distant semantic cues. By combining with cross-map interleaved scan (CIS), local pixel-wise interaction with positional inductive bias can also be introduced to SSMX-Corr. The implicit low-level motion descriptor is proposed as a non-visual prompt to enhance tracking robustness, addressing the intermittent tip visibility problem. Extensive experiments on a dataset with motorized needle insertion in both phantom and tissue samples demonstrate that the proposed tracker outperforms other state-of-the-art trackers while ablation studies further highlight the effectiveness of each proposed tracking module.
@@ I INTRODUCTION
In various percutaneous intervention procedures, ultrasound (US)-guided needle insertion is commonly adopted in minimally invasive interventions, such as tissue biopsy, tumor ablation, regional anesthesia [1], etc. As a non-invasive, portable, safe, and cost-effective imaging modality [2], US provides real-time intraoperative imaging of the needle and tissue, thereby minimizing the risk of accidental injury to vessels or critical organs. Despite these advantages, US imaging has inherent limitations, including low resolution and susceptibility to noise and artifacts [1], which can obscure or distort the needle tip position. Furthermore, under the narrow US imaging plane [2], small structures, such as the needle tip, can disappear intermittently when they are obscured by anatomical structures or not co-planar with the US imaging plane. These challenges underscore the necessity for robust and accurate needle tracking to ensure successful insertion procedures under challenging environments.
Prior to the widespread adoption of learning-based needle trackers, traditional methods, such as the statistical filter [3] and Gabor filter [4], achieved somewhat satisfactory performance but were hindered by complex workflows and sensitivity to hyper-parameters, failing to address challenging environmental factors in US imaging. A method based on discriminative correlation filter (DCF) has been proposed in [5], but correlation filter-based methods can be susceptible to background distraction and image distortion [6]. Later, DCF [5] is integrated with an optical tracking system for higher accuracy [7], but it is constrained by the cumbersome deployment of the optical tracking system. Recently, deep learning methods based on convolutional neural network (CNN) and transformer [8] have gained popularity in tracking tasks [9, 10, 11, 12], including US needle tracking. Mwikirize et. al. proposed a US needle tracker with a two-step structure based on a fully convolutional network and region-based CNN [13]. A paradigm utilizing digital subtraction is proposed in [14], which augments tip features to enhance visibility prior to tracking. However, these two methods have non-end-to-end structures that require multiple steps to localize the needle tip, potentially affecting robustness and accuracy. Since the needle tip is often obscured or distorted under artifacts and noise, some other methods perform needle shaft segmentation before localizing the needle tip [15, 16]. Although segmentation mask of the needle shaft can provide useful cues on the axial position of needle tip, accumulative error and discrepancy can be accidentally introduced by this two-stage workflow, and again constrained by the disadvantages of non-end-to-end methods. Additional segmentation mask data is also required to train these segmentation models, leading to additional obstacles for the model deployment.
@@
In addition to the challenging environment with noise and artifacts, the intermittently visible needle tip poses another problem that hinders accurate needle tracking. This issue can be caused by deviation of US imaging plane or obstruction by anatomical structures [19]. Mwikirize et. al. proposed a single-shot needle tracker by integrating historical frames to enhance needle tip features, addressing scenarios where the tip is imperceptible or the shaft is invisible [20]. Although it integrates historical information, its feature pre-enhancement can unexpectedly shift the original latent features, causing potential inherent information loss. Integrating historical needle motion into visual tracking presents another approach, since motion information can serve as an effective non-visual prompt to prevent tracking failure when the needle tip is invisible. A motion prediction module is integrated with a visual tracker in [6], yet it is constrained by its explicit motion prediction that poses challenges on generalizability when encountering motion from unseen domains.
@@
Mamba [21, 22] has recently drawn considerable attention and is being applied in real-world tracking tasks [23, 24, 25]. Based on the structured state space models (SSMs) [26], Mamba has a computationally efficient long-range lossless modeling capability with its selective scan mechanism. This capability allows Mamba-based trackers by leveraging long-range information, such as aggregating a video-level template set [24] or integrating historical frames [23]. It should be noted that a transformer-based tracker can hardly achieve long-range modeling under similar model size and complexity since the transformer has quadratic time complexity and memory requirement. It also requires positional encoding that may not effectively capture lossless long-range dependencies compared to Mamba which is designed specifically for such tasks. Thus Mamba usually outperforms transformer-based methods under similar model size and complexity [21]. Since no Mamba-based tracker has yet been proposed for US needle tracking, developing an effective Mamba-based needle tracker under US imaging remains an open research area.
@@
To address the aforementioned challenges of noise, artifacts, and intermittent visibility in US needle tracking, in this paper, a Mamba-based US needle tracker MambaXCTrack utilizing SSM cross-correlation (SSMX-Corr) and implicit motion prompt is proposed. To the best of our knowledge, it is the first time a Mamba-based tracker has been adopted in US needle tracking. It is also the first time that cross-correlation is implemented with SSM. Leveraging the long-range modeling capability of SSMs, SSMX-Corr enables global search and long-range modeling of distant semantic features. When the tip feature is degraded by noise and artifacts, SSMX-Corr avoids tracking failure by implicitly learning distant semantic feature that potentially comes from visual cues like needle shaft, rather than by explicitly performing segmentation on the needle shaft like existing methods [15, 16]. By further integrating the proposed cross-map interleaved scan (CIS), SSMX-Corr enjoys global search without losing local pixel-wise interaction between search and template maps to keep positional inductive bias, while existing convolutional X-Corr models [9, 10] only consider local modeling. To address the intermittent visibility of the needle tip, the implicit low-level motion descriptor is introduced as a non-visual prompt in addition to visual features that can unpredictably become unreliable. Different from [6] that trains an external motion predictor to explicitly predict the future motion, the proposed workflow preprocesses motion to obtain the low-level motion descriptor, which is then implicitly integrated with visual features. This implicit low-level motion integration introduces image-agnostic raw motion and ensures an end-to-end network to enhance training stability and tracking robustness. Extensive evaluations on a dataset of motorized needle insertions in both phantom and animal tissue demonstrate MambaXCTrack’s superior performance compared to state-of-the-art methods. The main contributions are fourfold:
@@
  • SSMX-Corr improves existing cross-correlation with SSM by globally searching and long-range modeling distant semantic features, thus learning potential distant visual cues. This represents the first effort to adopt Mamba effectively for US needle tracking.
@@
  • Implicit low-level motion descriptor is adopted to provide a non-visual prompt, thus leveraging motion information to address the challenge of intermittent needle visibility to achieve robust and consistent tracking.
  • The proposed tracker achieves SOTA performance on both phantom and tissue experiments. Further ablation studies show the effectiveness of the proposed modules.
@@ II-A MambaXCTrack
The overview of MambaXCTrack is shown in Fig. 1. The template map $Z$ and the search map $X$ first go through the ResNet [27] backbone, from which the embedded features $z\in\mathbb{R}^{H_{z}\times W_{z}\times C}$ and $x\in\mathbb{R}^{H_{x}\times W_{x}\times C}$ ($C=128$) are obtained through a cascaded linear projection. In the following Mamba head, after going through cascaded layer normalization, linear projection (in the dimension of $2C$), convolution (kernel size $3\times 3$, stride 1), and SiLU activation [28], $x$, $z$, and the low-level motion descriptor $m$ are transmitted to SSMX-Corr. For each Mamba head, $x$ is serially passed to the next head for a better feature representation, yet $z$ and $m$ are concurrently transmitted by each head to avoid error accumulation on the original template and raw motion.
@@ II-C SSMX-Corr: SSM is Convolution with Global Search
where * denotes the convolution. Since convolutional X-Corr is a local operation, it only models local pixel-wise interaction yet fails to consider interrelationship between distant features. When local features from the small area of needle tip is degraded by severe noise and artifacts, X-Corr can lost tracking easily without seeking hints from distant semantic features. Searching globally for potential semantic visual cues then turns out to be a promising paradigm, inspired by segmentation-based needle tracker [15, 16], which segments needle shaft to obtain needle’s axial direction.
@@
where $\textbf{A}\in\mathbb{R}^{N\times N}$, $\textbf{B}\in\mathbb{R}^{N\times 1}$, and $\textbf{C}\in\mathbb{R}^{1\times N}$. It maps the input sequence $u(t)\in\mathbb{R}$ to output $y(t)\in\mathbb{R}$ through latent states $h(t)\in\mathbb{R}^{N\times 1}$. To discretize Eq. 2 for deep learning integration, $\Delta$ is introduced as the step size at timescale. Equation (2) is then discretized as
@@
where the discritized matrices $\bar{\textbf{A}}$ and $\bar{\textbf{B}}$ are given by $=exp({\Delta}\cdot\textbf{A})$ A ) and $=({\Delta}\cdot\textbf{A})^{-1}(exp({\Delta}\cdot\textbf{A})-I)\cdot{\Delta}\textbf{B}$ B. During the discretization, the selective scan mechanism [21] is adopted as a parameterization of SSM in an input-dependent manner, where the parameters $\bar{\textbf{B}},\textbf{C},\Delta$ are derived from the input sequence.
By reformulating Eq. 3, the closed form of the output $y_{k}$ at every time step $k$ can be derived as
@@
Then $y$ can be formulated as the result of a convolution
$y=\bar{\mathcal{K}}*u,$ (5)
where $\bar{\mathcal{K}}$ is the SSM kernel given by
$\begin{split}\bar{\mathcal{K}}&=(\textbf{C}\bar{\textbf{A}}^{j}\bar{\textbf{B}})_{j\in[L]}\in\mathbb{R}^{L}\\ &=(\textbf{C}\bar{\textbf{B}},\textbf{C}\bar{\textbf{A}}\bar{\textbf{B}},...,\textbf{C}\bar{\textbf{A}}^{L-1}\bar{\textbf{B}}).\end{split}$ (6)
@@
which is a convolution between the concatenated map from scanned aggregation $CIS(z,x_{i},m)$ and the kernel $\bar{\mathcal{K}}$, while $\bar{\mathcal{K}}$ comes from the parameterization of input, i.e., $z$, $x$, and $m$. Thus SSMX-Corr can be regarded as a self-convolution of $z$, $x$, and $m$, yet enjoys SSM’s long-range modeling capabilities. With SSMX-Corr, tip tracking can benefit from global search between $x$ and $z$. Distant semantic features from the needle shaft can be learned to guide tracking. The sequence modeling diagram in SSMX-Corr also facilitates better motion-vision information fusion than existing convolutional X-Corr.
@@ II-D Implicit Low-level Motion Descriptor
There exist some US needle trackers [6] that improve tracking performance by motion information. However, its explicit motion integration by a motion predictor makes it a non-end-to-end model, posing potential harm to its generalizability and training stability. The implicit low-level motion descriptor is introduced in this paper for robust tracking when the needle tip is intermittently invisible. Given a set of historical bounding boxes $B=\{\beta_{1},\beta_{2},...,\beta_{t}\}$ ($t$ is the time index), where $\beta_{t}=(w_{t},h_{t},cx_{t},cy_{t})$ is the bounding box defined by its width $w_{t}$, height $h_{t}$, and coordinate of the top-left corner $(cx_{t},cy_{t})$, the low-level motion descriptor $m_{t}$ is constructed using the local displacement, given by $m_{t}=(cx_{t}-cx_{t-1},cy_{t}-cy_{t-1})=(\Delta cx_{t},\Delta cy_{t})$. This low-level local displacement smooths out the distraction from absolute information at the picture level, benefiting the model inferencing and generalizability. The width and height are removed to avoid overfitting since target size is not informative enough for needle tip tracking. To store the historical motion during tracking, a motion queue at a length $T$ is built with a FIFO (first-in-first-out) strategy. Noted that for queue storage, the motion sequence is not sampled by time to keep the original motion step. The sequence of $m$ is then formulated as $\{m_{1},m_{2},...,m_{T}\}$ in the size of $T\times 2$, which is concatenated with $z$ and $x$ in SSMX-Corr for implicit information fusion.

### v2

@@ Abstract
Ultrasound (US)-guided needle insertion is widely employed in percutaneous interventions. However, providing feedback on the needle tip position via US imaging presents challenges due to noise, artifacts, and the thin imaging plane of US, which degrades needle features and leads to intermittent tip visibility. In this paper, a Mamba-based US needle tracker MambaXCTrack utilizing structured state space models cross-correlation (SSMX-Corr) and implicit motion prompt is proposed, which is the first application of Mamba in US needle tracking. The SSMX-Corr enhances cross-correlation by long-range modeling and global searching of distant semantic features between template and search maps, benefiting the tracking under noise and artifacts by implicitly learning potential distant semantic cues. By combining with cross-map interleaved scan (CIS), local pixel-wise interaction with positional inductive bias can also be introduced to SSMX-Corr. The implicit low-level motion descriptor is proposed as a non-visual prompt to enhance tracking robustness, addressing the intermittent tip visibility problem. Extensive experiments on a dataset with motorized needle insertion in both phantom and tissue samples demonstrate that the proposed tracker outperforms other state-of-the-art trackers while ablation studies further highlight the effectiveness of each proposed tracking module.
Index Terms:
AI-Based Methods, Deep Learning Methods, Computer Vision for Medical Robotics, Visual Tracking
@@ I INTRODUCTION
In various percutaneous intervention procedures, ultrasound (US)-guided needle insertion is commonly adopted in minimally invasive interventions, such as tissue biopsy, tumor ablation, regional anesthesia [1], etc. As a non-invasive, portable, safe, and cost-effective imaging modality [2], US provides real-time intraoperative imaging of the needle and tissue, thereby minimizing the risk of accidental injury to vessels or critical organs. Despite these advantages, US imaging has an inherent limitation of susceptibility to noise and artifacts [1], which can obscure or distort the needle tip position. Furthermore, under the narrow US imaging plane, small structures, such as the needle tip, can disappear intermittently when they are obscured by anatomical structures or not co-planar with the US imaging plane [2]. These challenges underscore the necessity for robust and accurate needle tracking to ensure successful insertion procedures under challenging environments.
Prior to the widespread adoption of learning-based needle trackers, traditional methods, such as the statistical filter [3] and Gabor filter [4], achieved somewhat satisfactory performance but were hindered by complex workflows and sensitivity to hyper-parameters, failing to address challenging environmental factors in US imaging. A method based on a discriminative correlation filter (DCF) has been proposed in [5], but correlation filter-based methods can be susceptible to background distraction and image distortion [6]. As a result, this DCF method has also been integrated with an optical tracking system for higher accuracy [7], but the deployment of an optical tracking system is a cumbersome constraint. Recently, deep learning methods based on convolutional neural networks (CNN) and transformers [8, 9] have gained popularity in tracking tasks [10, 11, 12], including US needle tracking [13, 14]. Mwikirize et. al. proposed a US needle tracker with a two-step structure based on a fully convolutional network and a region-based CNN [15]. A paradigm utilizing digital subtraction is proposed in [16], which augments tip features to enhance visibility prior to tracking. However, these two methods have non-end-to-end structures that require multiple steps to localize the needle tip, potentially affecting robustness and accuracy. In addition, since the needle tip is often obscured or distorted under artifacts and noise, some other methods perform needle shaft segmentation before localizing the needle tip [17, 18]. While a segmentation mask of the needle shaft can offer useful cues for determining the axial position of the needle tip, this two-stage workflow may inadvertently introduce accumulative errors and discrepancies. Extra data of segmentation mask is also required to train segmentation models, leading to additional obstacles for model deployment.
@@
In addition to the challenging environment with noise and artifacts, the intermittently visible needle tip poses another problem that hinders accurate needle tracking. This issue can be caused by deviation of the US imaging plane or obstruction by anatomical structures [22]. Mwikirize et. al. proposed a single-shot needle tracker by integrating historical frames to enhance needle tip features, addressing scenarios where the tip is imperceptible or the shaft is invisible [23]. Although it integrates historical information, its feature pre-enhancement can unexpectedly shift the original latent features, causing potential inherent information loss. Integrating historical needle motion into visual tracking presents another approach, since motion information can serve as an effective non-visual prompt to prevent tracking failure when the needle tip is invisible. A motion prediction module is integrated with a visual tracker in [6], yet it is constrained by its explicit motion prediction that poses challenges on generalizability when encountering motion from unseen domains.
@@
Mamba [24] has recently drawn considerable attention and is being applied in real-world tracking tasks [25, 26, 27]. Based on the structured state space models (SSMs) [28], Mamba has a computationally efficient long-range lossless modeling capability with its selective scan mechanism. Leveraging this, Mamba-based trackers can efficiently model long-range information, such as aggregating a video-level template set [26] or integrating historical frames [25]. It should be noted that a transformer-based tracker can hardly achieve long-range modeling under similar model size and complexity since the self-attention mechanism in transformers has quadratic time complexity and memory requirement with respect to the sequence length [8]. It also requires positional encoding that may not effectively capture lossless long-range dependencies compared to Mamba, which is designed specifically for such tasks. Thus, Mamba usually outperforms transformer-based methods under similar model size and complexity [24]. Since no Mamba-based tracker has yet been proposed for US needle tracking, developing one remains an open research area.
@@
To address the aforementioned challenges of noise, artifacts, and intermittent visibility in US needle tracking, in this work, MambaXCTrack, a Mamba-based US needle tracker utilizing SSM cross-correlation (SSMX-Corr) and an implicit motion prompt, is proposed. To the best of our knowledge, it is the first time a Mamba-based tracker has been adopted in US needle tracking. It is also the first time that cross-correlation is implemented with SSM. Leveraging the long-range modeling capability of SSMs, SSMX-Corr enables global search and long-range modeling of distant semantic features. When the needle tip feature is degraded by noise and artifacts, SSMX-Corr avoids tracking failure by implicitly learning distant semantic features that potentially come from visual cues like the needle shaft, rather than by explicitly performing segmentation on the needle shaft like existing methods [17, 18]. By further integrating the proposed cross-map interleaved scan (CIS), SSMX-Corr enjoys global search without losing local pixel-wise interaction between search and template maps to keep positional inductive bias, while existing convolutional X-Corr models [10, 29] only consider local modeling. To address the intermittent visibility of the needle tip, an implicit low-level motion descriptor is introduced as a non-visual prompt in addition to visual features that can unpredictably become unreliable. Different from [6] that trains an external motion predictor to explicitly predict the future motion, the proposed workflow preprocesses motion to obtain the low-level motion descriptor, which is then implicitly integrated with visual features. This implicit low-level motion integration introduces image-agnostic raw motion and ensures an end-to-end network to enhance training stability and tracking robustness. Extensive evaluations on a dataset of motorized needle insertions in both phantom and animal tissue demonstrate MambaXCTrack’s superior performance compared to state-of-the-art methods. The main contributions are fourfold:
@@
  • SSMX-Corr improves existing cross-correlation with SSM by globally searching and modeling long-range distant semantic features, thus learning potential distant visual cues. This represents the first effort to adopt Mamba effectively for US needle tracking.
@@
  • An implicit low-level motion descriptor is adopted to provide a non-visual prompt, thus leveraging motion information to address the challenge of intermittent needle visibility to achieve robust and consistent tracking.
  • The proposed tracker achieves state-of-the-art (SOTA) performance on both phantom and tissue experiments. Further ablation studies show the effectiveness of the proposed modules.
@@ II-A MambaXCTrack
The overview of MambaXCTrack is shown in Fig. 1. A ResNet-50 network [30] with 50 layers is adopted as the backbone. The template map $Z$ and the search map $X$ first go through the backbone, from which the embedded features $z\in\mathbb{R}^{H_{z}\times W_{z}\times C}$ and $x\in\mathbb{R}^{H_{x}\times W_{x}\times C}$ ($C=128$) are obtained through a cascaded linear projection. In the following Mamba head, after going through cascaded layer normalization, linear projection (in the dimension of $2C$), convolution (kernel size $3\times 3$, stride 1), and SiLU activation [31], $x$, $z$, and the low-level motion descriptor $m$ are transmitted to SSMX-Corr. For each Mamba head, $x$ is serially passed to the next head for a better feature representation, yet $z$ and $m$ are concurrently transmitted by each head to avoid error accumulation on the original template and raw motion.
@@ II-C SSMX-Corr: SSM is Convolution with Global Search
where * denotes the convolution. Since convolutional X-Corr is a local operation, it only models local pixel-wise interaction yet fails to consider interrelationship between distant features. When local features from the small needle tip area are degraded by severe noise and artifacts, X-Corr can lose tracking easily without seeking hints from distant semantic features. Searching globally for potential semantic visual cues then turns out to be a promising paradigm inspired by segmentation-based needle trackers [17, 18], which obtain the needle’s axial direction by segmentation.
@@
where $\textbf{A}\in\mathbb{R}^{N\times N}$ (state matrix), $\textbf{B}\in\mathbb{R}^{N\times 1}$ (input matrix), and $\textbf{C}\in\mathbb{R}^{1\times N}$ (output matrix). These two equations together map the input sequence $u(t)\in\mathbb{R}$ to output $y(t)\in\mathbb{R}$ through linear transformation of the latent states $h(t)\in\mathbb{R}^{N\times 1}$. To integrate the continuous system described in Eq. 2 into a digital system for deep learning integration, it needs discretization first. $\Delta$ is introduced as the step size at timescale. Eq. 2 is then discretized as
@@
where the discritized matrices $\bar{\textbf{A}}$ and $\bar{\textbf{B}}$ are given by $=exp({\Delta}\cdot\textbf{A})$ A ) and $=({\Delta}\cdot\textbf{A})^{-1}(exp({\Delta}\cdot\textbf{A})-I)\cdot{\Delta}\textbf{B}$ B. During the discretization, the selective scan mechanism [24] is adopted to learn input-dependent parameters $\textbf{B},\textbf{C},\Delta$ from input $x\in\mathbb{R}^{L\times C}$, where $L$ is the sequence length, $C$ is the dimension. Parameters $\textbf{B},\textbf{C},\Delta$ are given by $\textbf{B}=\text{Linear}(x)\in\mathbb{R}^{L\times N}$, $\textbf{C}=\text{Linear}(x)\in\mathbb{R}^{L\times N}$, $\Delta=\text{Softplus}(\widetilde{\Delta}+\text{Linear}(x))\in\mathbb{R}^{L\times C}$, where $\widetilde{\Delta}$ is a trainable parameter.
By reformulating Eq. 3 [32], the closed form of the output $y_{k}$ at every time step $k$ can be derived as
@@
Then $y$ can be formulated as the result of a convolution by extracting the coefficients into the SSM kernel $\bar{\mathcal{K}}$
$\begin{split}y&=\bar{\mathcal{K}}*u,\\ \bar{\mathcal{K}}&=(\textbf{C}\bar{\textbf{A}}^{j}\bar{\textbf{B}})_{j\in[L]}\in\mathbb{R}^{L}\\ &=(\textbf{C}\bar{\textbf{B}},\textbf{C}\bar{\textbf{A}}\bar{\textbf{B}},...,\textbf{C}\bar{\textbf{A}}^{L-1}\bar{\textbf{B}}).\end{split}$ (5)
@@
which is a convolution between the concatenated map from scanned aggregation $CIS(z,x_{i},m)$ and the kernel $\bar{\mathcal{K}}$. Note that $\bar{\mathcal{K}}$ comes from the parameterization of input ($z$, $x$, $m$). Thus, SSMX-Corr can be regarded as a self-convolution of $z$, $x$, $m$, yet it enjoys SSM’s long-range modeling capabilities. With SSMX-Corr, tip tracking can benefit from global search between $x$ and $z$. Distant semantic features from the needle shaft can be learned to guide tracking. The sequence modeling diagram in SSMX-Corr also facilitates better motion-vision information fusion than existing convolutional X-Corr.
@@ II-D Implicit Low-level Motion Descriptor
There exist some US needle trackers [6] that improve tracking performance by incorporating motion information. However, explicit motion integration by including a motion predictor makes it a non-end-to-end model, posing potential harm to its generalizability and training stability. An implicit low-level motion descriptor is introduced in this paper for robust tracking when the needle tip is intermittently invisible. Given a set of historical bounding boxes $B=\{\beta_{1},\beta_{2},...,\beta_{t}\}$ ($t$ is the time index), where $\beta_{t}=(w_{t},h_{t},cx_{t},cy_{t})$ is the bounding box defined by its width $w_{t}$, height $h_{t}$, and coordinate of the top-left corner $(cx_{t},cy_{t})$, the low-level motion descriptor $m_{t}$ is constructed using the local displacement, given by $m_{t}=(cx_{t}-cx_{t-1},cy_{t}-cy_{t-1})=(\Delta cx_{t},\Delta cy_{t})$. The motion descriptor $m_{t}$ is in pixels units. This low-level local displacement smooths out the distraction from absolute information at the picture level, benefiting the model inferencing and generalizability. To store the historical motion during tracking, a motion queue at a length $T$ is built with a FIFO (first-in-first-out) strategy. Note that for queue storage, the motion sequence is not sampled by time to keep the original motion step. The sequence of $m$ is then formulated as $\{m_{1},m_{2},...,m_{T}\}$ in the size of $T\times 2$, which is concatenated with $z$ and $x$ in SSMX-Corr for implicit information fusion.